Let (V,+V,·V) and (W,+W,·W) be vector spaces and define

V ⊕ W = {(v,w) : v ∈ V and w ∈ W}. Prove that

(a) V ⊕ W is a vector space, under componentwise operations.

(b) ifS={(v,0):v∈V}andS′ ={(0,w):w∈ W}, then S and S′ are subspaces of V ⊕ W.
Answer #1

@lv, +V;V) and (W, TW, ow) are too verton spaus NA W = {(6,W) | VE V & WEWE he have two prove that Vow is a verton Spar underSiskeur 40 ca P + P2 = (V , ed) + (va, wd) unes (here & (VI+V2, W, tud) EVO ) Pita f rtW (Vitva EV & with 6W row am veeton sp& Pitle= ( ) + (Va, wa). & vitvd, witwa) (va tvi , wat wa) & (datud ) + (Viti) & atli satisfies and commufective propenty. 1plis additine invense of p. in Vow. Hence for all pfrow 9 additive muense apfw such that pap= Orow. where p = (r, w) 8 pl= C(&TB) P = (t BJC v , w) Epf row) = (2+B)V, (2-4PW) alfasualan = (QV+BV, aw tow) = QV,2w) + (pv, pos) = 2(V, W) + B (V, W) dptd al sit so) fs (when S = 1 o) , a anu scalar Q & B (s,tsal) es (where s (0,01)] - ß cay sealan Sea (0,wal 0 d( $175x) = 4